OpenStudy (anonymous):

p + 11 = –13 A. –24 B. –2 C. 2 D. 24

OpenStudy (uri):

p=-13-11 p=-24

OpenStudy (anonymous):

is positive+positive=negitive?

sammixboo (sammixboo):

A

OpenStudy (uri):

That happens in multiplication,bro.

OpenStudy (anonymous):

Not supposed to give out the answeres

sammixboo (sammixboo):

No postitive + positive = negative

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

does positive-negitive=positve

sammixboo (sammixboo):

p + 11 = -13 p + 11 - 11 = -13 - 11 p = -13 + -11 p = -24